# Benchmarking Differentially Private Tabular Data Synthesis

### v1

@@ D.5. Hyperparameters for Different Synthesis Algorithms
Preprocessing Method Marginal Size ML Efficiency Query Error Fidelity Error
@@
Raw $51.05e+$ 5 $84.24e+$ 8 $82.88e+$ 8 $0.76$ $0.72$ $0.54$ $0.001$ $0.001$ $0.001$ $0.003$ $0.001$ $0.001$
Numerical preprocessing $26.48e+$ 2 $35.05e+$ 3 $61.14e+$ 6 $0.76$ $0.71$ $0.53$ $0.002$ $0.002$ $0.003$ $0.006$ $0.002$ $0.003$
Categorical Preprocessing $51.05e+$ 5 $84.24e+$ 8 $82.31e+$ 8 $0.76$ $0.72$ $0.54$ $0.001$ $0.001$ $0.001$ $0.003$ $0.001$ $0.001$
Full Preprocessing $26.48e+$ 2 $35.05e+$ 3 $33.14e+$ 3 $0.76$ $0.71$ $0.53$ $0.002$ $0.002$ $0.003$ $0.006$ $0.002$ $0.003$
@@
ML efficiency $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
PrivSyn selection $0.67$ $0.65$ $0.68$ $0.67$ $0.80$ $0.75$- -- -- -- -- -- -- -- --
@@
ML efficiency $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
@@
Appendix E Supplementary Experiment Results
E.1. More Results of Algorithm Comparison
We have provided some important metric results of algorithm utility in Section 9.1. Here we list some other detailed evaluations such as AUC, accuracy for machine learning efficiency and running time for time efficiency. Notably, these results do not necessarily influence our conclusion, thus we present them in the appendix.
E.2. Detailed Results of Preprocessing Influence
@@ E.3. Detailed Results of Reconstruction Experiment
The detailed results of different preprocessing methods (used to plot Figure 7 and Figure 8) are shown in Table 20 andTable 21. Here, there are some “-” in the table. This is because PrivSyn tends to select as many marginals as possible, which will form large cliques. This will cause the size of the graphical model to be too large, requiring extremely large memory.

### v2

@@ E.2. Detailed Results of Preprocessing Influence
Preprocessing Method Marginal Size ML Efficacy Query Error Fidelity Error
@@
Raw $1.05e+5$ $4.24e+8$ $2.88e+8$ $0.76$ $0.72$ $0.54$ $0.001$ $0.001$ $0.001$ $0.003$ $0.001$ $0.001$
Numerical preprocessing $6.48e+2$ $5.05e+3$ $1.14e+6$ $0.76$ $0.71$ $0.53$ $0.002$ $0.002$ $0.003$ $0.006$ $0.002$ $0.003$
Categorical Preprocessing $1.05e+5$ $4.24e+8$ $2.31e+8$ $0.76$ $0.72$ $0.54$ $0.001$ $0.001$ $0.001$ $0.003$ $0.001$ $0.001$
Full Preprocessing $6.48e+2$ $5.05e+3$ $3.14e+3$ $0.76$ $0.71$ $0.53$ $0.002$ $0.002$ $0.003$ $0.006$ $0.002$ $0.003$
@@
ML Efficacy $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
PrivSyn selection $0.67$ $0.65$ $0.68$ $0.67$ $0.80$ $0.75$ - - - - - - - - -
@@
ML Efficacy $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$ $\varepsilon=0.2$ $\varepsilon=1$ $\varepsilon=5$
@@ E.3. Detailed Results of Reconstruction Experiment
The detailed results of different preprocessing methods (used to plot Figure 8) are shown in Table 25 and Table 26. Here, there are some “-” in the table. This is because PrivSyn tends to select as many marginals as possible, which will form large cliques. This will cause the size of the graphical model to be too large, requiring extremely large memory.
E.4. Discussion for Cramer’s V measure
We further explain why we do not include Cramer’s V measure in our evaluation metrics with some experimental results. In Table 23, a key finding is that even though some methods vary largely in fidelity error, their evaluations of Cramer’s V measure error can be close. For example, GEM and RAP++ both have a result of 0.1 on Cramer’s V measure error, but their fidelity errors actually differ. This result supports our argument in Section 8.